OpenStudy (anonymous):

consider the function *Find the equation of the tangent to the curve when x=2

OpenStudy (anonymous):

\[f(x)=(x+b)^2*(2x-1)\]

OpenStudy (anonymous):

above equation is the subjected function of the queation..! help me!

hartnn (hartnn):

do u know how to find slope of tangent from f(x) ??

hartnn (hartnn):

NO?

hartnn (hartnn):

slope of tangents at x=2 is given by derivative of f(x) at x=2 does this ring the bell ?

hartnn (hartnn):

tell me what u get f'(x) as..... ?

OpenStudy (anonymous):

its \[f'(x)=6x^2+8xb+2b^2-2x-2b\]

hartnn (hartnn):

the slope of tangent at x= 2 is given by f'(2) so just put x=2 there, and find slope.

OpenStudy (anonymous):

so \[y-f(x)=f'(x)(x-2)\] seems to be right?

hartnn (hartnn):

the point will be (2,f(2))

hartnn (hartnn):

\(y-f(2)=f'(2)(x-2)\)

hartnn (hartnn):

and slope = f'(2)

hartnn (hartnn):

ok?

OpenStudy (anonymous):

thank you it seems much easirer!

hartnn (hartnn):

yes, its easy :) welcome ^_^

OpenStudy (anonymous):

then how can i find the value of b in exact form if the straight line with equation y=4x-2 intersects y=f(x) at the minimum turning point? what would be the easiest step?